SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

   
Prev Next


Date: January 1, 2008
Shift Taker: Chris West
-----------
General Features:

Masked out fiber 3-10 C.

(theta,phi)
Warm crystals in hitmap          (52,63), (57,63), (56,64)
Warm crystals in 0.01 GeV hitmap (32,64), (17,43), (30,45), (32,46)
Warm crystals in 0.1 GeV hitmap  (32,64), (30,45), (32,46)
Warm crystals in 1 GeV hitmap      

Cool crystal  in hitmap          
Cool crystal  in 0.01 GeV hitmap 
Cool crystal  in 0.1 GeV hitmap  (29,85)
Cool crystal  in 1 GeV hitmap    (3,43)

77785-77787: OK. 
77788: Beam related noisy channel at (1,43)
77789: OK.
77790: High backgrounds.
77791: OK.
77792: 1e-3 Out of Synch in 4-6.
77793: 1e-3 Out of Synch in 4-6 (with exactly zero other damages)
77794-77797:
77798: Large background spikes.
77799: Large background spkies. ~3e-4 OOO, OOS in 4-10
77800: Second peak in digi timing at t=6600 ns. 4e-4 OOO, OOS in 4-10
77801-77803: OK.
77804: 4e-4 OOO, OOS in 4-10
77805-77808: OK.
77809:4e-4 OOS in 1-6.
77810:  OK.
77811: Second peak in digi timing at t=6600 ns.



Crystal Spectra:
NTR.

Lightpulser:
27 dead channels, 0 channels out of time. Updated on: 30-December-07 09:05 PST
Big second blob in logE vs time plot.
channel vs rom oot plot is a mess.


 

  Links

  • Shift Report Index
  • Back to the Shifter instructions